OpenStudy (anonymous):

Find the total work done by gravity on a steel ball weighing 4oz as it rolls down a helical track parameterized by r(t) = with 0<=t<=8. I don't get it.. Find the total work done by gravity on a steel ball weighing 4oz as it rolls down a helical track parameterized by r(t) = with 0<=t<=8. I don't get it.. @Mathematics

OpenStudy (anonymous):

what's your r(t) lol

OpenStudy (anonymous):

\[r(t)=<\cos \pi t,\sin \pi t, 8-t>\]

OpenStudy (anonymous):

well w=fd correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and you know xi+yj+zk=r(t)?

OpenStudy (anonymous):

basically you're going to need to find the distance and the force of the ball

OpenStudy (anonymous):

i believe

OpenStudy (anonymous):

basically i believe this is an arclength problem in which it wants you to find also the force of the ball